OpenStudy (anonymous):

How would I change this into slope–intercept form? 2x + 3y = -6

OpenStudy (amistre64):

by solving for y

OpenStudy (ash2326):

We have \[2x+3y=-6\] Slope intercept is in th form of \[y=mx+c\] now let's bring x to the other side \[3y=-6-2x\] divide both sides by 3 \[y=-\frac{2}{3}x-2\] slope is \(-\frac{2}{3}\) y- intercept= -2

OpenStudy (anonymous):

Omgosh thank ya'll so much!